OpenStudy (anonymous):

Explain your answer Can the values on both sides of the equals sign be different?

OpenStudy (anonymous):

then what would the equal sign mean?

OpenStudy (anonymous):

No. When there is an equal sign, both sides of the equation must be equal in value.

OpenStudy (turingtest):

C'mon guys look at \[x+7=x-5\to7=-5\]which means the answer has no real solution so yes, they can be different, but really this is written as \[7\neq-5\]I guess the answer is s semantic issue... Can they be different? Yes, but the statement is untrue.

OpenStudy (turingtest):

vague question in my opinion...

OpenStudy (anonymous):

No. Equal means they are the same, if they are not the same then they are not equal. If you come up with an answer like that then the statement is untrue.